OpenStudy (anonymous):

There are new houses being built in a neighborhood. Last month, of them were sold. This month, of the remaining houses were sold. How many houses are left to be sold?

OpenStudy (anonymous):

There are new houses being built in a neighborhood. Last month, of them were sold. This month, of the remaining houses were sold. How many houses are left to be sold?

OpenStudy (anonymous):

How many were sold & how many of the remaining houses were sold? because the numbers aren't showing up

OpenStudy (anonymous):

75 new houses 1/5 were sold, this month 1/3

OpenStudy (anonymous):

There are 75 new houses being built in neighborhood Last month 1/5 were sold. This month 1/3 of remaining houses were sold

OpenStudy (anonymous):

How many houses are left to be sold

OpenStudy (anonymous):

There are 75 new houses being built in neighborhood Last month 1/5 were sold. This month 1/3 of remaining houses were sold . How many houses ar eleft to be sold?

OpenStudy (texaschic101):

75 new houses last month 1/5 were sold : 1/5(75) = 75/5 = 15 75 - 15 = 60 this month 1/3 of remaining were sold 1/3(60) = 60/3 = 20 60 - 20 = 40 there are 40 remaining houses to be sold

OpenStudy (texaschic101):

any questions ?